OpenStudy (anonymous):

limit x-->3- (x-2)/(x-3)

OpenStudy (anonymous):

\[\lim_{x\to3^-}\frac{x-2}{x-3}\] 1. at x=3, we have undefined. 2. when \(2<x<3\), numerator is positive and deno. is negative 3. Hence, the limit approaches \(\boxed{-\infty}\)